asiaimann:

The table below shows the number of different types of faucets sold over three months and the total selling price of the items. Number of Faucets Sold and Total Selling Price Kitchen Sink Bathroom Sink Shower Total Selling Price January 15 8 7 $2,830 February 17 12 8 $3,450 March 24 25 21 $6,680 The company makes a profit of 15% on every faucet sold. What is the profit on one shower faucet? $12.00 $13.50 $14.15 $18.00

Vocaloid:

this table is a bit confusing but I think I know how you're supposed to solve it for each month: (kitchen sink price) * (number of kitchen sinks) + (bathroom sink price) * (number of bathroom sinks) + (shower faucet price)(number of shower faucets) = total selling price for January: 15k + 8b + 7s = 2830 for February: 17k + 12b + 8s = 3450 for March: 24k + 25b + 21s = 6680 three equations, three unknowns. solve for s, then calculate 15% of s to see what the profit is in $